# PhreshPhish: A Real-World, High-Quality, Large-Scale Phishing Website Dataset and Benchmark

### v1

@@ 6 Conclusions
We introduced PhreshPhish, which to the best of our knowledge, constitutes the largest, highest-quality public corpus of phishing websites available. Existing phishing website datasets often suffer from low quality due to the inherent challenges of scraping phishing websites. To address this, we implemented a robust scraping pipeline and then systemetically post-processed the collected data to eliminate invalid and mislabeled samples (see Algorithm 1 in Appendix C). We also minimized train-test leakage (see Algorithm 2 in Appendix D), another concern with existing datasets. Most published work on phishing detection report excellent results – primarily due to use of unrealistically high base rates. As expected, we found that model performance degrades sharply as the base rate decreases to more realistic values. To enable more realistic evaluation of phishing models, we constructed $\sim 400$ benchmark datasets with real-world base rates ranging from 0.05% to 5%, while also increasing task difficulty and diversity (see Algorithm 3 in Appendix D). Our evaluation of common detection approaches on the benchmark datasets shows significant room for improvement. Given the importance of false positives on user experience, we adopt average precision and precision@recall=R as key performance metrics. We make our dataset and benchmarks publicly available on Hugging Face. The code used to scrape and process the dataset, construct the test and benchmark datasets, and to build and evaluate the baseline models is publicly available on Github https://github.com/phreshphish/phreshphish. Finally, to keep the dataset fresh, we plan to release versioned datasets and benchmarks periodically.
Appendix A PhreshPhish Datasheet
{tblr}
width = 1.1colspec = Q[valign=m,halign=l,2.2cm] Q[valign=m,halign=l,2.2cm] X[l], cell11 = r=2m, cell31 = r=4m, cell71 = r=2m, cell91 = r=2m, cell111 = r=3m, cell141 = r=3m, vlines = , hlines = , Motivation & Purpose To further anti-phishing research, particularly web-based phishing. One of the primary challenges of web-based phishing detection is that it is difficult to obtain high-quality, real world phishing datasets and we hope that this dataset and benchmark will help to spur more innovation and research in the field.
Creators OpenText, Inc. fully funded and created the dataset.
Composition Description Dataset is comprised of HTML webpage data, labels, and metadata in Parquet format.
[…]
    "sha256": "0b7244604505e864d5d836...",
    "url": "https://login-gemini-homep...",
    "label": "phishing",
    "target": "gemini",
    "date": "2025-03-02",
    "lang": "en",
    "lang_score": 0.893,
    "html": "<html>...</html>"
Count About 372k samples: 253k benign pages and 119k phishing pages.
Splits Train and test splits are provided as well as 404 benchmarks created from the test split.
Collection Process Acquisition The data was acquired by scraping the HTML contents located at a URL. URLs were sourced from various phishing and benign feeds.
Time Frame The data was scraped over the course of 8 months, between July 2024 and March 2025.
Data Processing Description The raw data underwent the following steps: (1) URL deduplication; (2) cleaning as described in Section 3; (3) Train/test split and benchmark creation as described in Section 4.
Code Availability All code is available on Github https://www.github.com/phreshphish/phreshphish.
Usage Considerations Intended Usage This dataset should be used to further anti-phishing research.
Cautions Because the URLs were sourced, in part, from customer use traces, some of the corresponding webpages may contain questionable content including pornographic and/or hateful speech.
Limitations Although effort was made to ensure the dataset is of high quality, due to the size it is impossible to fully validate that low-quality samples have been fully removed.
Distribution License The dataset is made available under the Creative Commons Attribution 4.0 International license.
Hosting The full dataset is publicly available on Hugging Face https://huggingface.co/datasets/phreshphish/phreshphish.
Maintenance We plan to periodically release updates to the dataset as new phishing URLs are discovered.
Appendix B Existing Datasets
We searched for phishing datasets on common dataset hosting sites such as Hugging Face and Kaggle. On each of these sites, we searched for the terms "phishing" and "phish". On Hugging Face, we restricted the search to datasets with more than 10K rows. On Kaggle, we restricted the search to datasets larger than 100 MB in size. Almost all datasets we found were either URL-only or related to email phishing, with only a few containing both raw HTML webpage and URL data.
Table 3: Phishing datasets on common hosting sites
Repository Total Feature type
URL only Email URL & HTML
Hugging Face 35 15 14 2
Kaggle 19 6 4 4
Harvard Dataverse 2 0 2 0
OpenML 4 2 1 0
UCI ML Repo 3 3 0 0
B.1 Data Quality
In this section, we provide a quality assessment of two datasets found on Kaggle: Aljofey [3] and Crawling2024 [15]. Our goal is to demonstrate that these exhibit substantial quality issues and therefore motivate the need for a new, high-quality dataset. We apply the same cleaning methodology used for constructing our own dataset, using rejection rates from manual inspection as a proxy for data quality. Specifically, we apply: (1) URL normalization and deduplication; (2) removal of samples with titles indicative of scrape failure (Table 4); (3) grouping based on TF-IDF features using LSH and manual inspection of the groups; and (4) grouping by title and manual inspection of the most common titles. To make the comparisons feasible across datasets, we set a budget of 60 LSH bins and 50 distinct titles: approximately one hour of work for a human annotator. High rejection rates among the inspected samples suggests that poor quality is not confined to edge cases but is pervasive. We summarize the results of our quality assessment in Table 5.
@@ B.1 Data Quality
Table 4: Keywords used for filtering bad titles. These keywords are used to automatically identify and remove low-quality samples from the dataset.
@@
Because a complete manual inspection is not feasible, we estimate a lower and upper bound for the amount of data that would remain were cleaning to continue to completion. The upper bound is a generous estimate of the number of samples that would remain after cleaning and is equal to the number in row labeled "After Stage 2" in Table 5. The lower bound is, we believe, a more realistic estimate of the number of samples that would remain after cleaning and is equal to the upper bound multiplied by the rejection rate observed in the manual inspection of the LSH bins and titles. The rejection rate is calculated as the number of samples removed divided by the total number of samples in the dataset.
Table 5: Two-stage cleaning pipeline results for three datasets. Stage 1 applies automated filtering; Stage 2 applies manual inspection using two grouping schemes: LSH and title. Note: Some examples were removed by multiple filters, so removal counts may overlap and not sum to exact totals.
Step Aljofey Crawling2024 PhreshPhish (ours)
Phish Benign Total Phish Benign Total Phish Benign Total
Stage 0: Initial Size
Initial Size 27,280 32,972 60,252 26,687 29,428 56,115 118,752 253,189 371,941
Stage 1: Automated Filtering
URL Duplicates 2,936 129 3,065 1,069 166 1,235 0 0 0
HTML Missing 1,843 1,103 2,946 1,427 103 1,530 0 0 0
Bad Title 12,715 5,018 17,733 6,322 665 6,987 0 0 0
After Stage 1 12,722 26,851 39,573 18,938 28,660 47,598 118,752 253,189 371,941
Stage 2: Manual Inspection
LSH Rejections 2,530 5,422 7,952 3,390 3,470 6,860 331 80 411
Title Rejections 794 183 977 555 47 602 447 27 474
After Stage 2 9,398 21,246 30,644 14,993 25,143 40,136 117,974 253,082 371,056
Total Removed 17,882 11,726 29,608 11,694 4,285 15,979 778 107 885
Rejection Rate (%) 65.5% 35.6% 49.1% 43.8% 14.6% 28.5% 0.65% 0.04% 0.24%
Aljofey, et al. [3]
This dataset contains 60k samples, with 27k labeled as phishing and 33k labeled as benign. We found that this dataset contained a large number of data points that could immediately be removed under our automated filtering process (the majority of which were due to having bad titles such as "404 Not Found"). After running our cleaning process, over 65% of the phishing samples in this dataset were removed, leaving only 9,398 phish samples remaining in the dataset. The benign samples were also heavily affected, with 35% of the benign samples being removed.
Assuming the remaining un-inspected samples are perfectly clean, the dataset would contain 9,398 phishing samples and 21,246 benign samples. Under more realistic assumptions and based on rejection rates observed, however, we estimate that the dataset would contain only around 3,242 phishing samples and 13,682 benign samples if cleaning were to continue to completion.
Crawling2024 [52]
This dataset contains 56k samples, with 27k labeled as phishing and 29k labeled as benign. Although this dataset performed better than Aljofey, there was still a significant amount of data that was removed after applying our cleaning process.
Assuming the remaining un-inspected samples are perfectly clean, the dataset would contain 14,993 phishing samples and 25,143 benign samples. Under more realistic assumptions and based on rejection rates observed, however, we estimate that the dataset would contain only around 8,400 phishing samples and 20,000 benign samples if cleaning were to continue to completion.
PhreshPhish (ours)
Our dataset contains 372k samples, with 119k labeled as phishing and 253k labeled as benign. Because we had already applied the cleaning process to our own dataset, the automated filtering stage didn’t result in any new samples being removed. After applying the manual inspection process, we found only a few hundred samples needing removal. This is expected and we do not claim that our dataset is free from all errors as an unlimited cleaning budget is not feasible. However, we believe the significantly lower rejection rates in our dataset compared to the other two datasets shows that our dataset is of much higher quality.
Appendix C Our Dataset
C.1 Cleaning details
Following the data collection phase, we conducted a dedicated cleaning stage to address residual noise, failure modes, and inconsistencies not resolved by the scraping pipeline alone. While our browser-based scraper was designed to mitigate many of the challenges associated with phishing data collection (e.g., cloaking, ephemeral content), scraping artifacts and systemic noise remained common, particularly among phishing pages. Representative examples of these artifacts are shown in Figure 4.
Figure 4: Multiple failure modes are frequently encountered when scraping phishing pages. Top Left: DNS resolution issues can prevent the page from being scraped. Top Right: Security takedown notices can prevent the page from being scraped. Bottom Left: URL shortening services will sometimes block access to the intended page. Bottom Right: Redirects to the legitimate target brand page can occur resulting in mislabeled data.
@@ C.1 Cleaning details
Our approach to dataset cleaning, described in Algorithm 1, is based on the idea of using a combination of two different grouping strategies to quickly identify and remove pages that are likely to correspond to a failure mode. Pages are first vectorized using TF-IDF features (line 4). These vectors are then projected into binary hash bins using random projection to form coarse similarity groups (lines 5 and 9). Pages are also grouped by their HTML <title> tags (lines 7 and 10), which is helpful for identifying pages that are likely to be derived from the same template or boilerplate code. Within each group, a single representative (prototype) page is selected and presented for manual inspection (lines 18 and 19). The human annotator is then asked to determine whether the page reflects a failure mode or otherwise invalid content (line 20). If the page represents a scrape failure, it and all of its nearest neighbors (under cosine similarity in the TF-IDF space) are removed from the dataset (line 21). Groups are prioritized by size, ensuring that the cleaning budget is spent efficiently (lines 16 and 17).
@@
4:$X\leftarrow\texttt{TFIDF}(\mathcal{D})$triangleright create TF-IDF features
5:$B\leftarrow\big{(}X\cdot\mathcal{N}(0,1)\big{)}\geq 0$triangleright bin via random projection
@@
9:$G_{B}\leftarrow\texttt{GroupBy}(B,X)$triangleright group by LSH bins
10:$G_{T}\leftarrow\texttt{GroupBy}(T,X)$triangleright group by titles
@@
15:     $i\leftarrow 0$triangleright stop when budget exhausted
@@
[…]
@@
27:return $D^{\prime}\leftarrow X\setminus R$
C.2 Handling PII
@@
C.3 Risks and limitations
@@
C.4 Summary of selected dataset facets
Table 6: Summary statistics for selected URL features.
@@ C.4 Summary of selected dataset facets
URL length Num characters in the URL 14 8 83 44 40 58 57 78 2,544 1,199
Domain length Num characters in domain 1 1 6 6 7 9 10 12 60 43
Path length Num characters in the path 3 32 1 19 3 32 16 52 2,073 981
Subdomain length Num characters in subdomain 0 0 4 3 11 3 17 3 92 63
Figure 6: Distribution of URL characteristics.
Table 7: Summary statistics for selected HTML features.
@@
HTML length <html> tag lengths 42 23 69,66 65,320.75 38,400 180,243 126,522 409,249 39,544,635 57,863,340
Head tag length Sum of all <head> tag lengths 13 13 1,267 9,435 6,218 40,352 17902 118,630 25,817,150 30,385,165
Image tag length Sum of all <img> tag lengths 6 6 84 116 148 155 311 258 25,817,150 1,454,356
Script tag length Sum of all <script> tag lengths 17 17 92 108 165 204 311 597 2,257,807 13,280,918
Figure 7: Distribution of HTML characteristics.
Figure 8: Distribution of languages in the dataset.
Figure 9: Top phishing targets.
Appendix D Test and Benchmark Sets
Figure 5 provides a temporal view of the dataset after the cleaning process, aggregated at a daily granularity. The volume of phishing data collected each day depended primarily on the number of phishing pages reported by the feeds, as well as occasional local infrastructure issues. In contrast, the benign data collection shows greater variability, partly because benign URLs are more readily available – allowing a large number to be collected in a single day – and also because benign pages tend to be less dynamic and remain fresh for longer periods than phishing pages. This accounts for the periods during which no benign data was collected.
Algorithm 2 outlines the main procedure used to split the dataset into training and test sets. A test split fraction, $p=0.7$ was used. Lines 4--7 create a temporal split, performed in a stratified manner to preserve class proportions across both sets. However, due to differences in data collection between phishing and benign, the temporal separation between train and test occur at different times for the two classes as shown in Figure 5.
@@ Appendix D Test and Benchmark Sets
Algorithm 3 outlines the procedure for creating the benchmark datasets from the test set. Lines 5--6 describe the process of enhancing test dataset diversity. The benign part of the test set is augmented with legitimate websites corresponding to the top target phishing brands (determined from the training phish set). For the phishing test data, pairwise comparisons are performed, and one sample from each pair of highly similar instances is removed. This pruning step eliminates highly similar samples, increasing the diversity of the phishing dataset.
Lines 9--13 describe the difficulty filter, where a fraction of randomly selected “easy” points, defined by highly confident correct scores predicted by a model, are discarded. We use $\delta=0.15$, and $k=10\%$.
Lines 16–24 describe the rate adjustment procedure, where for each of five base rates, the number of phishing points ($numPhish$) and instances ($numInstances$) are determined, and subsequently for each base rate $numInstances$ samples of size $numPhish$ are sampled without replacement from the reduced phishing test set, obtained from the previous step (difficulty filter).
@@
6:$\mathcal{D}_{train}\leftarrow\{\mathcal{D}_{1:k}\mid k=\lfloor(1-p)*|\mathcal{D}|\rfloor\}$
7:$\mathcal{D^{\prime}}_{test}\leftarrow\mathcal{D}\setminus\mathcal{D}_{train}$
@@
20:return $\mathcal{D}_{train},\mathcal{D}_{test}$
Figure 5: The train and test datasets are temporally disjoint for each class. Note: the benign test set shows a spike in the number of pages collected in March 2025; these are pages that were collected by scraping the Google search results of top targeted brands.
@@
5:$\mathcal{D}_{new}\leftarrow\mathcal{D}_{test}\cup\mathcal{D}_{targetBenign}$
6:$\mathcal{D}_{new}\leftarrow\mathcal{D}_{new}\setminus\mathcal{D}_{similarHTML}$
@@
9:$m\leftarrow\text{fitModel}(\mathcal{D}_{train})$
10:$\text{score}\leftarrow m.\text{predict}(\mathcal{D}_{new})$
11:$\mathcal{D}_{easy}\leftarrow\{x\mid\text{score}(x)>(1-\delta)\wedge x\in P\}\cup\{x\mid\text{score}(x)<\delta\wedge x\in B\}$
12:$\mathcal{D}_{discard}\leftarrow\text{sample}(k,\mathcal{D}_{easy})$
13:$\mathcal{D}_{new}\leftarrow\mathcal{D}_{new}\setminus\mathcal{D}_{discard}$
@@
16:$\mathcal{D}_{benign}\leftarrow\{x\mid x\in\mathcal{D}_{new}\wedge x\in B\}$
17:$\mathcal{D}_{phish}\leftarrow\{x\mid x\in\mathcal{D}_{new}\wedge x\in P\}$
$\mathbf{\mathcal{B}}\leftarrow\emptyset$
19:for $b\in\{0.05,0.1,0.5,1,5\}$ do
20:     $numPhish\leftarrow\lfloor\frac{b\times|\mathcal{D}_{benign}|}{100-b}\rfloor$
21:     $numInstances\leftarrow getInstances(b)$
22:     $\mathcal{D^{\prime}}_{phish}\leftarrow\text{sample}(\mathcal{D}_{phish},numPhish,numInstances,replace=\mathtt{False})$
23:     $\mathbf{\mathcal{B}}\leftarrow\mathbf{\mathcal{B}}\cup\{\mathcal{D}_{benign},\mathcal{D^{\prime}}_{phish}\}$
24:end for
25:return $\mathcal{B}$
Appendix E Baseline Experiments
E.1 Setup, Hyperparameters and Training Details

### v2

@@ 6. Conclusions
We introduced PhreshPhish, which to the best of our knowledge, constitutes the largest, highest-quality public corpus of phishing websites available. Existing phishing website datasets often suffer from low quality due to the inherent challenges of scraping phishing websites. To address this, we implemented a robust scraping pipeline and then systemetically post-processed the collected data to eliminate invalid and mislabeled samples (see Algorithm 1 in Appendix A). We also minimized train-test leakage (see Algorithm 2 in Appendix B), another concern with existing datasets. Most published work on phishing detection report excellent results – primarily due to use of unrealistically high base rates. As expected, we found that model performance degrades sharply as the base rate decreases to more realistic values. To enable more realistic evaluation of phishing models, we constructed 975 benchmark datasets with real-world base rates ranging from 0.05% to 5%, while also increasing task difficulty and diversity (see Algorithm 3 in Appendix B). Our evaluation of common detection approaches on the benchmark datasets shows significant room for improvement. Given the importance of false positives on user experience, we adopt average precision and precision@recall=R as key performance metrics. We make our dataset and benchmarks publicly available on Hugging Face. The code used to scrape and process the dataset, construct the test and benchmark datasets, and to build and evaluate the baseline models is publicly available on Github https://github.com/phreshphish/phreshphish. Finally, to keep the dataset fresh, we plan to release versioned datasets and benchmarks periodically.
Appendix A Our Dataset
Figure 4. The train and test datasets are temporally disjoint for each class.
Our data collection system runs continuously, collecting URLs from multiple feeds, and scraping the corresponding HTML content using a real browser to ensure high fidelity. Figure 4 provides a temporal view of the dataset after the cleaning process, aggregated at a daily granularity. The volume of phishing data collected each day depended primarily on the number of phishing pages reported by the feeds, as well as occasional local infrastructure issues. In contrast, the benign data collection shows greater variability, partly because benign URLs are more readily available – allowing a large number to be collected in a single day – and also because benign pages tend to be less dynamic and remain fresh for longer periods than phishing pages. This accounts for the periods during which no benign data was collected.
A.1. Cleaning details
Cleaning consists of two stages: (1) a fully automated stage that removes pages with invalid HTML content and other obvious artifacts such as bad titles as shown in Figure 5; and (2) a human-in-the-loop stage that identifies and removes pages likely to correspond to failure modes. While our scraping system was designed to mitigate many of the challenges associated with phishing data collection described in Section 3.1, scrape failures are still common and must be removed. Representative examples of common failure modes are shown in Figure 5.
@@ A.1. Cleaning details
Table 5. Keywords used for filtering bad page titles. These keywords are used to automatically identify and remove low-quality samples from the dataset.
@@
During the manual inspection stage, pages are first vectorized using TF-IDF features (line 4). These vectors are then projected into binary hash bins using random projection to form coarse similarity groups (lines 5 and 9). Pages are also grouped by their HTML <title> tags (lines 7 and 10), which is helpful for identifying pages that are likely to be derived from the same template or boilerplate code. Within each group, a single representative (prototype) page is selected and presented for manual inspection (lines 18 and 19). The human annotator is then asked to determine whether the page reflects a failure mode or otherwise invalid content (line 20). If the page represents a scrape failure, it and all of its nearest neighbors (under cosine similarity in the TF-IDF space) are removed from the dataset (line 21). Groups are prioritized by size, ensuring that the cleaning budget is spent efficiently (lines 16 and 17).
@@
[…]
@@
[…]
@@
15:  $i\leftarrow 0$ $\triangleright$ stop when budget exhausted
@@
17:   for $g\in\texttt{Sorted}(G,\texttt{key}=|g|,\texttt{descending=True})$ do
@@
27:return $\mathcal{D}^{\prime}\leftarrow X\setminus R$
A.2. Handling PII
@@
Figure 5. Multiple failure modes are frequently encountered when scraping phishing pages. (5(a)) DNS resolution issues can prevent the page from being scraped. (5(b)) Security takedown notices can prevent the page from being scraped. (5(c)) URL shortening services will sometimes block access to the intended page. (5(d)) Redirects to the legitimate target brand page can occur resulting in mislabeled data.
A.3. Risks and Limitations
@@
Figure 6. Summary statistics for selected URL features.
@@
URL length Num characters in the URL 12 5 31 42 39 55 58 74 25,523 1,414
Domain length Num characters in domain 4 4 10 10 11 13 15 16 64 46
Path length Num characters in the path 0 0 1 17 1 29 14 48 6,597 1,414
Subdomain length Num characters in subdomain 0 0 3 3 9 3 17 3 92 81
Figure 7. Distribution of URL characteristics with 99th percentile outliers removed.
Figure 8. Summary statistics for selected HTML features.
@@
HTML length <html> tag lengths 28 23 8,271 90,570 34,672 249,954 117,473 545,994 76,259,693 45,877,378
Head tag length Sum of all <head> tag lengths 0 0 1,323 11,168 5,375 49,107 21,087 150,520 16,327,295 32,183,109
Image tag length Sum of all <img> tag lengths 0 0 0 895 397 4,070 3,154 12,673 13,288,010 44,236,577
Script tag length Sum of all <script> tag lengths 0 0 438 8,573 3,397 28,939 24,532 115,009 16,327,282 21,785,672
Figure 9. Distribution of HTML characteristics with 99th percentile outliers removed.
Figure 10. Dataset statistics: (a) top languages as determined by the langdetect library, (b) top phishing targets, and (c) top domains (both phishing and benign).
Appendix B Test and Benchmark Sets
Algorithm 2 outlines the main procedure used to split the dataset into training and test sets. A test split fraction, $p=0.3$ was used. Lines 5--7 create a temporal split, performed in a stratified manner to preserve class proportions across both sets.
@@ Appendix B Test and Benchmark Sets
Algorithm 3 outlines the procedure for creating the benchmark datasets from the test set.
Lines 5--6 describe the process of enhancing test dataset diversity. The benign part of the test set is augmented with legitimate websites corresponding to the top target phishing brands (determined from the training phish set). For the phishing test data, pairwise comparisons are performed, and one sample from each pair of highly similar instances is removed. This pruning step eliminates highly similar samples, increasing the overall diversity of test set.
Lines 9--13 describe the difficulty filter, where a fraction of randomly selected “easy” phishing points, defined by highly confident correct average scores predicted by three baseline models (Linear, FFN, and GTE), are discarded. We use $\delta=0.15$, and $k=10\%$.
Lines 16–25 describe the rate adjustment procedure, where for each of five base rates, the number of phishing points ($n_{p}$) and instances ($n_{i}$) are determined, and subsequently for each base rate $n_{i}$ samples of size $n_{p}$ are sampled without replacement from the reduced phishing test set.
@@
[…]
@@
20:
21:return $\mathcal{D}_{train},\mathcal{D}_{test}$
@@
[…]
@@
[…]
@@
[…]
18:
[…]
20:for $b\in\{0.05,0.1,0.5,1,5\}$ do
21:  $n_{p}\leftarrow\lfloor\frac{b\cdot|\mathcal{D}_{benign}|}{100-b}\rfloor$ $\triangleright$ num phish needed
22:  $n_{i}\leftarrow\texttt{GetInstances}(b)$ $\triangleright$ num instances needed
23:  $\mathcal{D^{\prime}}_{phish}\leftarrow\texttt{RandomSample}(\mathcal{D}_{phish},n_{p},n_{i},\texttt{replace=False})$
24:  $\mathbf{\mathcal{B}}\leftarrow\mathbf{\mathcal{B}}\cup\{\mathcal{D}_{benign},\mathcal{D^{\prime}}_{phish}\}$
25:end for
26:
27:return $\mathcal{B}$
B.1. Dependence of Model Performance on Base Rate
The performance of a classifier degrades with decrease in the base rate (fraction of the postive class) of a test data set. Note that by model performance we mean precision (or variants) here. While recall (true positive rate) and false positive rate (FPR) are not affected by base rate (assuming a fixed, trained model), precision is sensitive to it. In fact, one can write precision as a function of recall, FPR, and base rate as:
$\text{Prec}=\frac{R\cdot r}{R\cdot r+\text{FPR}(1-r)}$
where $R$ is recall (or true positive rate), FPR is false positive rate and $r$ is base rate. It is easy to check that $\frac{\partial\text{Prec}}{\partial r}$ is always positive, implying a monotonic, increasing relationship between precision and base rate; so as base rate drops, precision tends to go down. Further, the above partial derivative quantifies the sensitivity of precision to base rate.
Appendix C Baseline Experiments
C.1. Setup, Hyperparameters and Training Details